40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 12924     Entry time: Mon Apr 3 17:09:47 2017
Author: gautam 
Type: Update 
Category: CDS 
Subject: C1PSL burt-restored 

When I came in this morning, Steve had re-locked the PMC and IMC - but I could see a ~1Hz intensity fluctuation on the PMC REFL video monitor. I unlocked the PMC and tried to re-lock it, but couldn't using the usual prescription of turning the servo gain down and moving the DC bias slider around. I checked the status of the slow machines - all were responding to pings and could be telnet'ed into, so that didn't seem to be the problem. In the past, this sort of behaviour was characteristic of the infamous "sticky slider" problem - so I simply burt-restored c1psl using a snapshot from 29 March, after which I could easily re-lock the PMC. The transmitted light level looked normal on the scope on the PSL table, and the PMC REFL video monitor also look normal now.

ELOG V3.1.3-